# FedAdamW: A Communication-Efficient Optimizer with Convergence and Generalization Guarantees for Federated Large Models

### v2

@@ Abstract
AdamW has become one of the most effective optimizers for training large-scale models. We have also observed its effectiveness in the context of federated learning (FL). However, directly applying AdamW in federated learning settings poses significant challenges: (1) due to data heterogeneity, AdamW often yields high variance in the second-moment estimate $\boldsymbol{v}$; (2) the local overfitting of AdamW may cause client drift; and (3) Reinitializing moment estimates ($\boldsymbol{v}$, $\boldsymbol{m}$) at each round slows down convergence. To address these challenges, we propose the first Federated AdamW algorithm, called FedAdamW, for training and fine-tuning various large models. FedAdamW aligns local updates with the global update using both a local correction mechanism and decoupled weight decay to mitigate local overfitting. FedAdamW efficiently aggregates the mean of the second-moment estimates to reduce their variance and reinitialize them. Theoretically, we prove that FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$ without heterogeneity assumption, where $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds. We also employ PAC-Bayesian generalization analysis to explain the effectiveness of decoupled weight decay in local training. Empirically, we validate the effectiveness of FedAdamW on language and vision Transformer models. Compared to several baselines, FedAdamW significantly reduces communication rounds and improves test accuracy. The code is available in https://github.com/junkangLiu0/FedAdamW.
@@ Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (mcmahan2017communication) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (bian2025fedalt; wang2025adaptive; NEURIPS2024_a11e42a3; Bian_2025_ICCV; 10546478). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning (li2025multi; li2023ultrare; li2024survey; an2022numerical; an2024inspired; an2024robust; liu2023single; liu2025fedadamw; liu2025dp; liu2025fedmuon).
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (radford2018improving), RoBERTa (liu2019roberta), and Vision Transformers (ViT) (dosovitskiy2020image)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(bottou2010large) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (zhang2024transformers; liuimproving), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (zhang2024transformers). In contrast, AdamW (loshchilov2017fixing), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (vaswani2017attention; liu2019roberta), offering faster convergence and improved generalization, compared to Adam (kingma2014adam) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (mcmahan2017communication) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
@@
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (liu2019roberta), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
@@
Empirical results on ViT and LLMs confirm that FedAdamW improves test accuracy and reduces communication overhead compared to strong FL baselines.
@@
  • Empirical importance of AdamW and challenges in FL. We empirically demonstrate the effectiveness of the AdamW optimizer in federated settings, particularly for training Transformer models. Our analysis reveals three key challenges when applying AdamW in FL: (1) high variance in second-moment estimation, (2) local overfitting leading to client drift, and (3) inefficient convergence due to repeated initialization of optimizer states.
@@
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (li2018federated) introduces a proximal term to restrict local updates; SCAFFOLD (karimireddy2020scaffold) applies control variates to correct client drift; and FedCM (xu2021fedcm) leverages client momentum to stabilize updates. FedNSAM (liu2025consistency) analyzed the consistency between local and global flatness, FedBCGD  (liu2024fedbcgd) proposed a communication-efficient accelerated block coordinate gradient method. FedSWA  (liuimproving) further improved generalization under highly heterogeneous data via stochastic weight averaging.
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (duchi2011adaptive), Adadelta (zeiler2012adadelta), Adam (kingma2014adam), AMSGrad (reddi2019convergence), and AdamW (loshchilov2017fixing). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (liu2019roberta; zhang2024transformers; DBLP:journals/tnn/ZhangHLDCW25; zhang2025molebridge; DBLP:conf/aaai/Zhang0LXCCW25; DBLP:conf/aaai/ZhangLXCW024; ding2025traffic; zeng2025FSDrive; zeng2025janusvln; zhang2025fedagc; ENCODER; FineCIR; OFFSET; zhao2024balf; zhao2023benchmark; liao2025convex; liao2024globalpointer; chen2025improving; dang2023ticoserec; dang2025efficient; li2025dual; yang2025distillation; yang2023adaptive; qian2024bird; qian2025vovtrack; limva; li2024savsr; zhou2023fastpillars; pillarhist; zhou2024information; zhou2024lidarptq; ouyang2025multi; wei2025multiple; wei2025comprehensive; ouyang2025graph; ouyang2025dynamic; ouyang2025beaver).
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (reddi2020adaptive) incorporates server-side adaptivity using Adam and Yogi. FAFED (wu2023faster) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (chen2020toward) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, sun2023efficient proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
@@
Despite the widespread use of AdamW (loshchilov2017fixing; vaswani2017attention) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
@@
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (chen2020toward).
@@
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (collobert2004large; zhang2024adam) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
@@
These assumptions are standard in federated adaptive optimization literature (fan2024locally; sun2023efficient).
@@
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (shalev2014understanding). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
@@
Figure 6: Training loss curves on CIFAR-100 using ResNet-18 and ViT-Tiny under Dir-0.1, Dir-0.6. FedAdamW converges the fastest.
@@
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (krizhevsky2009learning), and Tiny ImageNet (le2015tiny). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2 (socher2013recursive), QQP (socher2013recursive). To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (hsu2019measuring), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (he2016deep) as a representative convolutional neural network (CNN), (ii) Swin Transformer (liu2021swin) and ViT-Tiny (dosovitskiy2020image) for Vision Transformers, and (iii) RoBERTa-Base (liu2019roberta) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (mcmahan2017communication), SCAFFOLD (karimireddy2020scaffold), FedCM (xu2021fedcm), FedAdam (reddi2020adaptive), FedLADA (sun2023efficient), Local Adam and Local AdamW.
@@
textbfQuestions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
Table 2: Comparison of test accuracy and training loss for Swin Transformer under Dir-0.1 with 100 communication rounds (100 clients, 5% participation, batch size 16, $K=50$).
@@
Table 3: Test accuracy (%) using RoBERTa-Base with LoRA across seven GLUE tasks over 100 communication rounds (mean ± std).
@@
Table 11 and Figure 6 present the test accuracy and training loss on CIFAR-100 using ResNet-18. FedAdamW achieves the best performance under both Dir-0.6 and Dir-0.1 settings, reaching a top accuracy of 66.12% and 63.01%, respectively. It also attains the lowest training loss (0.122 and 0.480), demonstrating faster and more stable convergence. Compared to other adaptive baselines such as FedAdam and Local AdamW, FedAdamW shows superior generalization under data heterogeneity, confirming its effectiveness in CNNs (Q3).
@@
Table 5: Impact of $\alpha$ using ViT-Tiny on CIFAR-100 (Dir-0.1).
$\alpha$ 0.00 0.25 0.50 0.75 1.00
@@
Impact of Aggregation Strategy. Table 7 shows that our strategy, Agg-mean-$\boldsymbol{v}$, achieves the best balance between accuracy and communication cost. While Agg-$\boldsymbol{v}$ improves performance by reducing variance, full aggregation (Agg-$\boldsymbol{vm}$) introduces excessive communication with marginal gains. In contrast, Agg-mean-$\boldsymbol{v}$ attains similar benefits with only $\mathcal{O}(B)$ communication—where $B$ is the number of blocks—demonstrating its scalability and effectiveness in stabilizing updates under strict communication constraints.
@@ Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (chen2023symbolic) or Lion (chen2023symbolic).

### v3

@@ Abstract
AdamW has become one of the most effective optimizers for training large-scale models. We have also observed its effectiveness in the context of federated learning (FL). However, directly applying AdamW in federated learning settings poses significant challenges: (1) due to data heterogeneity, AdamW often yields high variance in the second-moment estimate $\boldsymbol{v}$; (2) the local overfitting of AdamW may cause client drift; and (3) reinitializing moment estimates ($\boldsymbol{v}$, $\boldsymbol{m}$) at each round slows down convergence. To address these challenges, we propose the first Federated AdamW algorithm, called FedAdamW, for training and fine-tuning various large models. FedAdamW aligns local updates with the global update using both a local correction mechanism and decoupled weight decay to mitigate local overfitting. FedAdamW efficiently aggregates the mean of the second-moment estimates to reduce their variance and reinitialize them. Theoretically, we prove that FedAdamW achieves a linear speedup convergence rate of $\mathcal{O}(\sqrt{(L\Delta\sigma_{l}^{2})/(SKR\epsilon^{2})}+(L\Delta)/R)$ without heterogeneity assumption, where $S$ is the number of participating clients per round, $K$ is the number of local iterations, and $R$ is the total number of communication rounds. We also employ PAC-Bayesian generalization analysis to explain the effectiveness of decoupled weight decay in local training. Empirically, we validate the effectiveness of FedAdamW on language and vision Transformer models. Compared to several baselines, FedAdamW significantly reduces communication rounds and improves test accuracy.
Codehttps://github.com/junkangLiu0/FedAdamW
Extended versionhttps://arxiv.org/pdf/2510.27486
@@ Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (McMahan et al. 2017) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (Bian et al. 2025a; Wang et al. 2024; Bian et al. 2025b, 2024). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning (Li et al. 2025a, 2023; An et al. 2022, 2024a, 2024b; Liu et al. 2023, 2025d, 2025e, 2025c).
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (Radford et al. 2018), RoBERTa (Liu et al. 2019), and Vision Transformers (ViT) (Dosovitskiy et al. 2020)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(Bottou 2010) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (Zhang et al. 2024c; Liu et al. 2025a), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (Zhang et al. 2024c). In contrast, AdamW (Loshchilov et al. 2017), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (Vaswani et al. 2017; Liu et al. 2019), offering faster convergence and improved generalization, compared to Adam (Kingma and Ba 2014) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (McMahan et al. 2017) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
@@
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (Liu et al. 2019), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
@@
  • Empirical importance of AdamW and challenges in FL. We empirically demonstrate the effectiveness of the AdamW optimizer in federated settings, particularly for training Transformer models. Our analysis reveals three key challenges when applying AdamW in FL.
@@
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (Li et al. 2020) introduces a proximal term to restrict local updates; SCAFFOLD (Karimireddy et al. 2020) applies control variates to correct client drift; and FedCM (Xu et al. 2021) leverages client momentum to stabilize updates. FedNSAM (Liu et al. 2025b) analyzed the consistency between local and global flatness, FedBCGD  (Liu et al. 2024) proposed a communication-efficient accelerated block coordinate gradient method. FedSWA  (Liu et al. 2025a) further improved generalization under highly heterogeneous data via stochastic weight averaging.
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (Duchi et al. 2011), Adam (Kingma and Ba 2014), AMSGrad (Reddi et al. 2019), and AdamW (Loshchilov et al. 2017). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (Liu et al. 2019; Zhang et al. 2024c; Ouyang et al. 2025; Qian et al. 2024; Li et al. 2025b; Yang et al. 2025; Zhang et al. 2025, 2024b, 2024a; Wei et al. 2025; Zhou et al. 2023, 2024)
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (Reddi et al. 2020) incorporates server-side adaptivity using Adam and Yogi. FAFED (Wu et al. 2023) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (Chen et al. 2020) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, Sun et al. (2023) proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
@@
Despite the widespread use of AdamW (Loshchilov et al. 2017; Vaswani et al. 2017) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
@@
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (Chen et al. 2020).
@@
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (Collobert 2004; Zhang et al. 2024d) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
@@
These assumptions are standard in federated adaptive optimization literature (Fan et al. 2024; Sun et al. 2023).
@@
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (Shalev-Shwartz and Ben-David 2014). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
@@
Figure 6: Training loss curves on CIFAR-100 using ResNet-18 and ViT-Tiny under Dir-0.1, Dir-0.6.
@@
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (Krizhevsky et al. 2009), and Tiny ImageNet (Le and Yang 2015). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2, QQP. To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (Hsu et al. 2019), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (He et al. 2016) as a representative convolutional neural network (CNN), (ii) Swin Transformer (Liu et al. 2021) and ViT-Tiny (Dosovitskiy et al. 2020) for Vision Transformers, and (iii) RoBERTa-Base (Liu et al. 2019) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (McMahan et al. 2017), SCAFFOLD (Karimireddy et al. 2020), FedCM (Xu et al. 2021), FedAdam (Reddi et al. 2020), FedLADA (Sun et al. 2023), Local Adam and Local AdamW.
@@
Questions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
@@
Table 2: Comparison of test accuracy and training loss for Swin Transformer under Dir-0.1 with 100 communication rounds(100 clients, 5% participation, batch size 16, K=50K\!=\!50).
@@
Table 3: Test accuracy (%) using RoBERTa-Base with LoRA across seven GLUE tasks over 100 communication rounds.
@@
Table 1 and Figure 6 present the test accuracy and training loss on CIFAR-100 using ResNet-18. FedAdamW achieves the best performance under both Dir-0.6 and Dir-0.1 settings, reaching a top accuracy of 66.12% and 63.01%, respectively. It also attains the lowest training loss (0.122 and 0.480), demonstrating faster and more stable convergence. Compared to other adaptive baselines such as FedAdam, FedAdamW shows superior generalization under data heterogeneity, confirming its effectiveness in CNNs (Q3).
@@
$\alpha$ (Dir-0.1) 0.00 0.25 0.50 0.75 1.00
@@
Table 5: Impact of $\alpha$ using ViT-Tiny on CIFAR-100.
@@
Impact of Aggregation Strategy. Table 7 shows that our strategy, Agg-mean-$\boldsymbol{v}$, achieves the best balance between accuracy and communication cost. While Agg-$\boldsymbol{v}$ improves performance by reducing variance, full aggregation (Agg-$\boldsymbol{vm}$) introduces excessive communication with marginal gains. In contrast, Agg-mean-$\boldsymbol{v}$ attains similar benefits with only $\mathcal{O}(B)$ communication, where $B$ is the number of blocks, demonstrating its scalability and effectiveness in stabilizing updates.
@@ Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (Chen et al. 2023) or Lion (Chen et al. 2023).